Opened 2 years ago

Last modified 4 months ago

#583 new flight processing

ASTC15_28, flight day 168/2015, Albacete

Reported by: dac Owned by:
Priority: immediate Milestone: 2015 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by dac)

Data location: ~arsf/arsf_data/2015/flight_data/spain/ASTC15_28-2015_168_Albacete_summer_school

Data arrived from ARSF via SATA disk on 09/07/2015

Scientific objective: Flight for ATSC summer school

Priority: Unknown

PI: Alasdair Mac Arthur

CAMF Code N028

Sensors:

  • Fenix (requested)
  • LiDAR (flown, not requested)
  • Owl (flown, not requested)

Change History (24)

comment:1 Changed 2 years ago by dac

  • Description modified (diff)

comment:2 Changed 2 years ago by asm

Last edited 2 years ago by asm (previous) (diff)

comment:3 Changed 2 years ago by asm

Last edited 2 years ago by asm (previous) (diff)

comment:4 Changed 2 years ago by asm

Last edited 2 years ago by asm (previous) (diff)

comment:5 Changed 2 years ago by gej

Started Navigation Processing

GPS data doesn't have L2C so have to specify manually at -0.25 L2 cycles
Elevation mask set to 8.0 degrees
Kar maximum distance for dual frequency set to 120km due to distance of teru basestation

Using two base stations,
Alba data:
Lat: 38 58 40.51042
long: -1 51 23.03504
Ell Height: 751.655m
teru data:
Lat: 40 21 01.80637
Long: -1 07 27.46418
Ell. Height: 956.120m

Accuracy is within acceptable levels (below 0.35)
Combined Seperation: East -0.05 North -0.10

Navigation Processing Finished

Last edited 2 years ago by gej (previous) (diff)

comment:6 Changed 2 years ago by gej

Fenix SCT values are

Flightline 1 couldn't be processed due to not having correct flight times (start and end identical).

use_nav false in config file due to issues with sync times. So had to use QGis for intial line ups of flightlines (using google sattelite data).

Flightline FENIX
1
2 2.60
3 3.36
4 4.96
5 3.40
6 11.33
7 1.30
Last edited 2 years ago by gej (previous) (diff)

comment:7 Changed 2 years ago by gej

Hyperspectral Ready for Delivery Check

comment:8 Changed 2 years ago by dac

Fenix Processing

Waiting on updated calibration to process fenix data with.

comment:9 Changed 22 months ago by gej

Fenix Processing

Reprocessed with updated calibration file. Creating new Delivery.

comment:10 Changed 22 months ago by gej

Fenix Processing
Created delivery, Ready for DC.

comment:11 Changed 22 months ago by dac

Fenix Delivery Check

Starting delivery check

comment:12 Changed 22 months ago by dac

Fenix Delivery Check

Initial comments

  • In readme change "Flightline 1 not processable". Line 1 wasn't a proper line, in the logsheet the file numbers start at 2. Don't need to say line 1 wasn't processed, but do need to say why the numbering starts at 2, something like: "The instrument was turned on before data collection started, therefore numbering starts at 2"
  • Not quite sure what happened with the apl commands in the config file - most of the names were wrong (3b for 1b files), have fixed these so check_apl_cmd now runs.

Still need to check spectra so will do this on Monday.

comment:13 Changed 22 months ago by gej

Fenix Processing

Have changed comment in Readme to the above.

comment:14 Changed 22 months ago by dac

Fenix Delivery Check - continued

  • Spectra look fine, good match to 6S simulated spectra.
  • Different spectral VNIR binning was used for 2 - 5 and 6 - 7. This isn't noted anyware in the ticket. Should be added to Readme.

Once Readme has been updated can be delivered when updated data quality report is available.

Last edited 22 months ago by dac (previous) (diff)

comment:15 Changed 22 months ago by gej

Fenix Processing

Updated Readme to mention that VNIR binning is different between flightlines 2-5 and 6-7.

Waiting on updated Data Quality Report.

comment:16 Changed 22 months ago by dac

Fenix

Updated Readme is better. Have started zipping mapped files,

comment:17 Changed 22 months ago by dac

  • Description modified (diff)

Fenix

Copied updated data quality report. Delivered data to PI via FTP (arsf12).

comment:18 Changed 22 months ago by dac

Archiving

Started archiving for distribution by NEODC

comment:19 Changed 22 months ago by dac

Archiving

Data uploaded to NEODC.

comment:20 Changed 20 months ago by dac

Archiving

Data now available at NEODC - leaving ticket open for atmospheric correction component.

comment:21 Changed 17 months ago by dac

Re-generating IGM files for mapping atmospherically corrected data. Will remove once complete.

comment:22 Changed 4 months ago by dac

The zipped version of f168023b_mapped looks like it has been corrupted, the write date is later than the rest.
Will regenerate and upload to NEODC.

comment:23 Changed 4 months ago by dac

Regenerated zipped file. Version at NEODC is OK.

comment:24 Changed 4 months ago by dac

Atmospherically corrected hyperspectral data using an empirical line correction and ATCOR have been copied to the server under: ~arsf/arsf_data/2015/misc/ATSC15_28_atmospheric_correction.

The files corrected using ATCOR are being uploaded to NEODC as part of the summer school.

Note: See TracTickets for help on using tickets.